### v1

@@ 4.2 Waveform models
DiffWave. [44] is a diffusion probabilistic model designed for both conditional and unconditional waveform generation. We utilize the official implementation. [^8] https://github.com/lmnt-com/diffwave.
HiFiGAN. [45] is a widely known GAN-based waveform model that uses multiple transposed convolution blocks to progressively upsample and transform input mel-spectrograms into speech waveforms. The generator is optimized using multiple discriminator losses, a feature matching loss, and an L1 loss between the generated and ground truth mel-spectrograms. We utilize the HiFiGAN V1 architecture from the official implementation. [^9] https://github.com/jik876/hifi-gan.
Parallel WaveGAN. [46] is a lightweight vocoder model. It utilizes a non-autoregressive WaveNet [47] architecture combined with multi-resolution Short-Time Fourier Transform (STFT) loss and waveform adversarial loss. We used the official implementation. [^10] https://github.com/kan-bayashi/ParallelWaveGAN.
NSF HiFiGAN. [48] is similar to Parallel WaveGAN but explicitly incorporates a sine-based source signal as input to the generator. Additionally, it includes a noise branch that transforms random noise into an aperiodic signal. This aperiodic signal is combined with the generator’s periodic output for harmonic-plus-noise speech waveform generation. We used the official implementation. [^11] https://github.com/nii-yamagishilab/project-NN-Pytorch-scripts.
BigVGAN. [49] is a universal GAN-based vocoder that generalizes effectively across diverse scenarios, including unseen speakers, languages, and recording environments. By utilizing periodic activation functions and anti-aliased representations, BigVGAN introduces a beneficial inductive bias for speech synthesis. We used the official implementation [^12] https://github.com/NVIDIA/BigVGAN.
WaveGlow. [50] generates waveforms through a series of neural network-based invertible affine transformations conditioned on input mel-spectrograms. During training, the model parameters are optimized to whiten the ground-truth waveform as much as possible. We utilized the same toolkit with NSF HiFiGAN.
@@ 4.3 E2E and speech-language models with neural codecs
VALL-E, Multi-Scale Transformer, and Delay. VALL-E [2] predicts the first token of each frame using an autoregressive module, followed by a non-autoregressive prediction for the remaining tokens. Multi-Scale Transformer [51] employs a global Transformer for inter-frame modeling and a local Transformer for intra-frame modeling, maintaining full autoregression without approximation. In Delay [52], the multi-stream token sequences are processed using a “delay” interleave pattern, which enables approximate autoregressive prediction for both inter- and intra-frame modeling, achieving high efficiency. We utilized implementations of the three models in the ESPnet toolkit.4
MQTTS. [3] is designed to synthesize speech using real-world data from YouTube and podcasts. To address misalignments common in mel-spectrogram-based autoregressive models, it employs a multi-codebook vector quantization approach to improve both speech intelligibility and diversity. MQTTS aligns closely with the goals of this work, as we aim to develop a dataset that spans real-world data for both bona fide and spoofed speech. We used the official implementation. [^13] https://github.com/b04901014/MQTTS.
VITS. [53] is an E2E TTS model that combines a conditional variational autoencoder with stochastic duration prediction to generate waveforms from textual input. The model uses normalizing flow to learn latent representations from speech, while the stochastic duration predictor captures diverse speech prosody from text. For waveform generation, adversarial loss is employed to produce high-quality waveforms from the latent representations. We trained VITS using the ESPnet toolkit.4
@@ 4.4 Attack generation, partitioning, and protocols
Diverse combinations of acoustic and waveform models, alongside E2E and SpeechLM models, result in a total of 23 spoofing attacks. This approach is inspired by previous work, which demonstrated that both acoustic and waveform models impact the perceptual quality of synthesized speech [54]. Table 3 provides a detailed overview of the 23 spoofing attacks included in the SpoofCeleb dataset.
Data partitioning for SpoofCeleb required a more sophisticated approach compared to existing ASV or SDD datasets. An SDD dataset only needs to consider the binary bonafide or spoof label, while an ASV dataset focuses on speaker identities. SpoofCeleb, as a dataset for both SDD and SASV, must account for both bona fide/spoof labels and speaker identities simultaneously.
Speakers. For the speaker partitioning, we divide the 1,251 speakers in the bona fide data into three sets: 1,17111711,1711 , 171 for training, $4040$ for validation, and $4040$ for evaluation. This ensures that there are no overlapping speakers between any of the partitions.
@@
In the validation set, there are 6 attacks: A06, A07, and A11 to A14, combined with the bona fide data (A00). Attacks A06 and A07 represent known attacks from unknown speakers. Attacks A11 and A12 involve the same architecture as other attacks but differ in model training details. Specifically, A11 is fully trained from scratch using the TITW dataset, while in A02, the decoder was pre-trained. Similarly, A12 is fully trained from scratch on TITW, whereas A04 was pre-trained on LibriSpeechGigaSpeech [55], and the English subset of Multilingual LibriSpeech [33], and then fine-tuned on TITW. Attacks A13 and A14 serve as partially known attacks. In A13, the acoustic model (GradTTS) is known, but the waveform model (NSF HiFiGAN) is unknown. Similarly, in A14, the acoustic model (Matcha-TTS) is known, but the waveform model (HiFiGAN) is unknown.
Figure 2: Illustration of how the SpoofCeleb dataset is partitioned.
Table 2: Specification of the number of speech files and protocols. The number of trials equals to the number of speech files for SDD protocols.
@@
In the evaluation set, there are 9 attacks, A15 to A23. Attacks A15 and A16 involve known architectures but differ in configurations. For A15, the decoder is initialized with a pre-trained model, and the speaker embeddings are taken from target utterances, simulating a scenario where an attacker has access to the target speaker’s utterance pool. A16 was pre-trained using the same data composition as A04. Attacks A17 and A18 represent partially known attacks where the acoustic models are known, but the waveform models are not. Finally, A19 to A23 are fully unknown attacks, meaning no part of their models was encountered during training.
Figure 2 illustrates the three partitions of SpoofCeleb and Table 2 provides the statistics of each partition. In total, SpoofCeleb contains over 2.5 million speech samples.
@@
Table 3: Spoofing attacks of the SpoofCeleb database. There exist 23 different attacks stemming from 23 different TTS systems. : pre-trained, : decoder is pre-trained, : speaker embeddings from target utterances.
@@
Table 4: Quality and strength of 23 spoofing attacks included in SpoofCeleb dataset. SPF-EER (%) measures how hard the attack is to be rejected by a pre-trained ASV system. MCD, UTMOS, and DNSMOS demonstrate how noisy the attacks are and WER (%) measures the intelligibility.
@@ 5.1 SDD
Two E2E SDD models, RawNet2 [56] and AASIST [57], are used as the baseline systems. The RawNet2 model for SDD is an adapted version of RawNet2 originally designed for ASV. It features an input layer that processes raw waveforms directly and uses convolution-based residual blocks. Frame-level representations are aggregated, projected, and then passed through a binary classification head.
AASIST is one of the most widely adopted SDD models in recent literature. Like RawNet2, it includes an input layer that processes raw waveforms and utilizes convolution-based residual blocks. However, unlike RawNet2, AASIST incorporates graph attention network-based modules designed to capture spectral and temporal spoofing artifacts separately. It then employs heterogeneous stacking of graph attention layers to jointly model spectral and temporal information concurrently.
@@ 5.2 SASV
We employ three models as SASV baselines, all of which utilize the SKA-TDNN architecture [58]. These models are used to assess the impact of different training data and scenarios. SKA-TDNN is a convolution-based model with residual connections, incorporating dedicated modules and architectural design choices for multi-scale processing. It is an advanced version of the ECAPA-TDNN architecture [59].
Among the three SASV baseline models, the first model (“Conventional ASV”) is trained as a conventional ASV system using the VoxCeleb1&2 datasets, without considering spoof robustness. We utilize a pre-trained model from ESPnet-SPK [60]. The second model (“SASV trained on out-of-domain data”) is trained as an SASV model but uses out-of-domain data from the ASVspoof2019 logical access dataset [9]. We employ a pre-trained model from [61]. The third model (“SASV trained on SpoofCeleb”) is trained as an SASV model using the training set from SpoofCeleb.
@@ 6 Metrics
A diverse set of metrics is employed to evaluate the SpoofCeleb dataset, as well as the SDD and SASV models. To assess the quality of the speech samples and the strength of the attacks, we use SPF-EER, Mean Cepstral Distortion (MCD), UTMOS [62], DNSMOS [63], and Word Error Rate (WER), with WER evaluated using the OpenAI Whisper-Large model [64]. For evaluating the performance of SDD baselines, we use Equal Error Rate (EER) and the min Detection Cost Function (minDCF) [65]. To assess the SASV baselines, we adopt the recently proposed architecture-agnostic Detection Cost Function (min a-DCF) [66], along with Speaker Verification EER (SV-EER) and SPooF EER (SPF-EER). Table 5 outlines the trial types involved in SASV metrics; a-DCF includes all three trial types, while SV-EER and SPF-EER cover only a subset.
Table 5: Three metrics used for gauging the performance of SASV baselines. a-DCF measures the overall performance. SV-EER measures the ability to reject non-target speakers and SPF-EER measures the spoof-robustness. “+”: a system should accept, “-”: a system should reject.
@@
Table 6: Speech deepfake detection baseline performances.
@@
Table 7: Attack-wise performance of RawNet2 SDD baseline on the validation and evaluation sets. Performances reported using SPF-EER (%).
@@
Table 8: SASV baseline performances. SKA-TDNN [58] model architecture is employed. The three models are trained under different scenarios. “Conventional ASV” is trained with VoxCelebs1&2 [14, 67] and “SASV trained on out-of-domain data” is trained on ASVspoof2019 logical access [9].
@@ 7.1 Spoofing attacks
In the top row, the speech quality evaluations for A00 (bona fide speech) are provided as reference values. The results confirm that the spoofing attacks in SpoofCeleb are highly threatening, with most attacks achieving an SPF-EER over 20%. Additionally, the majority of attacks exhibit relatively minor degradation in UTMOS and DNSMOS, indicating the high quality of the synthesized speech samples. Furthermore, intelligibility, measured by WER, shows that for most attacks, there is no more than a 10% deterioration in performance.
@@ 7.2 SDD
Table 6 presents the results of four baseline SDD systems. We evaluate two SDD models, RawNet2 and AASIST, trained on two different datasets. The models trained on the ASVspoof2019 logical access dataset are used to assess the zero-shot performance on SpoofCeleb validation and evaluation SDD protocols. The other two models demonstrate the performance of systems trained on in-domain SpoofCeleb training data.
The zero-shot results in the top two rows indicate that existing SDD models not trained on in-the-wild data struggle to distinguish between spoofed samples and bonafide speech. In rows 3 and 4, we observe a significant improvement in performance when these models are trained using the SpoofCeleb training data, highlighting the importance of training SDD models on in-the-wild data. However, RawNet2’s result in row 3 is unexpected, as it shows better performance on the evaluation set than on the validation set while the evaluation set includes totally unknown attacks. To further investigate, we conduct an analysis of the attack-wise results.
@@
Through further analysis, we found that the inferior performance on attack A11 contributed to the validation set results being worse than those on the evaluation set. Interestingly, when comparing A11 and A15, attack A15 is more difficult to distinguish for a conventional ASV system that doesn’t account for spoofing, with SPF-EER values of 47.78% for A11 and 65.21% for A15. Both attacks originate from MQTTS; however, A11 was trained entirely from scratch, while A15 utilized a pre-trained decoder. Once an SASV system is trained on the SpoofCeleb training data, A11 becomes more challenging to detect. A deeper investigation into the reasons behind this phenomenon is left for future work.
@@ 7.3 SASV
Table 8 presents the performance of three SASV baselines on the SpoofCeleb validation and evaluation protocols. Min a-DCF assesses the overall performance, while SV-EER and SPF-EER evaluate the systems’ ability to reject bona fide and spoof non-target trials, respectively.
As expected, a conventional ASV system that does not account for spoof attacks, shown in the first row, fails to reject synthesized speech samples, with an a-DCF exceeding 0.49 on both the validation and evaluation sets. However, it performs well at rejecting bona fide non-target trials. The results in the second row show an improvement in a-DCF for the validation set, but even worse performance on the evaluation set. Both SV-EER and SPF-EER remain very high, indicating that the system trained for SASV with out-of-domain data struggles to reject both types of non-target trials. a-DCF of 0.9998 also signifies that the model fails to find an operating point where it can reject both types of non-target trials. Finally, when trained on the SpoofCeleb training data, the a-DCF on the evaluation set drops to its lowest value (0.2902), and both SV-EER and SPF-EER are more balanced compared to row 1, where the system was only capable of rejecting bonafide non-target trials.

### v2

@@ 4.2 Waveform models
DiffWave. [44] is a diffusion probabilistic model designed for both conditional and unconditional waveform generation. We used the official implementation. [^8] https://github.com/lmnt-com/diffwave.
HiFiGAN. [45] is a widely known GAN-based waveform model that uses multiple transposed convolution blocks to progressively upsample and transform input mel-spectrograms into speech waveforms. The generator is optimized using multiple discriminator losses, a feature matching loss, and L1 loss between the generated and ground truth mel-spectrograms. We used the HiFiGAN V1 architecture from the official implementation. [^9] https://github.com/jik876/hifi-gan.
Parallel WaveGAN. [46] is a lightweight vocoder model. It uses a non-autoregressive WaveNet [47] architecture combined with multi-resolution Short-Time Fourier Transform (STFT) loss and waveform adversarial loss. We used the official implementation. [^10] https://github.com/kan-bayashi/ParallelWaveGAN.
NSF HiFiGAN. [48] is similar to Parallel WaveGAN but explicitly incorporates a sine-based source signal as input to the generator. It also includes a noise branch that transforms random noise into an aperiodic signal. This aperiodic signal is combined with the generator’s periodic output for harmonic-plus-noise speech waveform generation. We used the official implementation. [^11] https://github.com/nii-yamagishilab/project-NN-Pytorch-scripts.
BigVGAN. [49] is a universal GAN-based vocoder that generalizes effectively across diverse scenarios, including unseen speakers, languages, and recording environments. By using periodic activation functions and anti-aliased representations, BigVGAN introduces a beneficial inductive bias for speech synthesis. We used the official implementation [^12] https://github.com/NVIDIA/BigVGAN.
WaveGlow. [50] generates waveforms through a series of neural network-based invertible affine transformations conditioned on input mel-spectrograms. During training, the model parameters are optimized to whiten the ground-truth waveform as much as possible. We used the same toolkit as with NSF HiFiGAN.
@@ 4.3 E2E and speech-language models with neural codecs
VALL-E, Multi-Scale Transformer, and Delay. VALL-E [2] predicts the first token of each frame using an autoregressive module, followed by a non-autoregressive prediction for the remaining tokens. Multi-Scale Transformer [51] uses a global Transformer for inter-frame modeling and a local Transformer for intra-frame modeling, maintaining full autoregression without approximation. In Delay [52], the multi-stream token sequences are processed using a “delay” interleave pattern, which enables approximate autoregressive prediction for both inter- and intra-frame modeling, achieving high efficiency. We used implementations of the three models in the ESPnet toolkit.4
MQTTS. [3] is designed to synthesize speech using real-world data from YouTube and podcasts. To address misalignments common in mel-spectrogram-based autoregressive models, it uses a multi-codebook vector quantization approach to improve both speech intelligibility and diversity. MQTTS aligns closely with the goals of this work, as we aim to develop a dataset that spans real-world data for both bona fide and spoofed speech. We used the official implementation. [^13] https://github.com/b04901014/MQTTS.
VITS. [53] is an E2E TTS model that combines a conditional VAE with stochastic duration prediction to generate waveforms from textual input. The model uses normalizing flow to learn latent representations from speech, while the stochastic duration predictor captures diverse speech prosody from text. For waveform generation, adversarial loss is used to produce high-quality waveforms from the latent representations. We trained VITS using the ESPnet toolkit.4
@@ 4.4 Attack generation, partitioning, and protocols
Diverse combinations of acoustic and waveform models, alongside E2E and SpeechLM models, result in a total of 23 spoofing attacks. This approach is inspired by previous research, which demonstrated that both acoustic and waveform models impact the perceptual quality of synthesized speech [54]. Table 3 provides a detailed overview of the 23 spoofing attacks included in SpoofCeleb.
Data partitioning for SpoofCeleb requires a more sophisticated approach compared to existing ASV or SDD datasets. An SDD dataset only requires the binary bona fide or spoof label, while an ASV dataset focuses on speaker identities. SpoofCeleb, as a dataset for both SDD and SASV, must account for both bona fide/spoof labels and speaker identities simultaneously.
Speakers. For the speaker partitioning, we divide the 1,251 speakers in the bona fide data into three sets: 1,17111711,1711 , 171 for training, $4040$ for validation, and $4040$ for evaluation. This ensures that there are no overlapping speakers between any of the sets.
@@
In the validation set, there are 6 attacks: A06, A07, and A11 to A14, combined with the bona fide data (A00). Attacks A06 and A07 represent known attacks from unknown speakers. Attacks A11 and A12 involve the same architecture as other attacks but differ in model training details. Specifically, A11 is fully trained from scratch using the TITW dataset, while in A02, the decoder was pre-trained. Similarly, A12 is fully trained from scratch on TITW, whereas A04 was pre-trained on LibriSpeechGigaSpeech [55] and the English subset of Multilingual LibriSpeech [33], then fine-tuned on TITW. Attacks A13 and A14 serve as partially known attacks. In A13, the acoustic model (GradTTS) is known, but the waveform model (NSF HiFiGAN) is unknown. Similarly, in A14, the acoustic model (Matcha-TTS) is known, but the waveform model (HiFiGAN) is unknown.
Figure 2: Illustration of how SpoofCeleb is partitioned.
Table 2: Number of speech files and protocols. Number of trials equals that of speech files for SDD protocols.
@@
In the evaluation set, there are 9 attacks, A15 to A23. Attacks A15 and A16 involve known architectures but differ in configurations. For A15, the decoder is initialized with a pre-trained model, and the speaker embeddings are taken from target utterances, simulating a scenario in which an attacker has access to the target speaker’s utterance pool. A16 was pre-trained using the same data composition as A04. Attacks A17 and A18 represent partially known attacks where the acoustic models are known, but the waveform models are not. Finally, A19 to A23 are fully unknown attacks, meaning no part of their models was encountered during training.
Figure 2 illustrates the three partitions of SpoofCeleb and Table 2 provides the statistics of each partition. In total, SpoofCeleb contains over 2.5 M speech samples.
@@
Table 3: Spoofing attacks of SpoofCeleb. There are 23 different attacks stemming from 23 different TTS systems. : pre-trained, : decoder is pre-trained, : speaker embeddings from target utterances.
@@
Table 4: Quality and strength of 23 spoofing attacks included in SpoofCeleb. SPF-EER (%) measures how hard it is to reject an attack by a pre-trained ASV system. MCD, UTMOS, and DNSMOS demonstrate how noisy the attacks are and WER (%) measures the intelligibility.
@@ 5.1 SDD
Two E2E SDD models, RawNet2 [56] and AASIST [57], are used as the baselines. The RawNet2 model for SDD is an adapted version of RawNet2 originally designed for ASV. It features an input layer that processes raw waveforms directly and uses convolution-based residual blocks. Frame-level representations are aggregated, projected, then passed through a binary classification head.
AASIST is one of the most widely used SDD models in recent literature. Like RawNet2, it includes an input layer that processes raw waveforms and uses convolution-based residual blocks. However, unlike RawNet2, AASIST incorporates graph attention network-based modules designed to capture spectral and temporal spoofing artifacts separately. It then uses heterogeneous stacking of graph attention layers to jointly model spectral and temporal information concurrently.
@@ 5.2 SASV
We employ three models as SASV baselines, all of which use the SKA-TDNN architecture [58]. These models are used to assess the impact of different training data and scenarios. SKA-TDNN is a convolution-based model with residual connections, incorporating dedicated modules and architectural design choices for multi-scale processing. It is an advanced version of the ECAPA-TDNN architecture [59].
Among the three SASV baselines, the first model (“Conventional ASV”) is trained as a conventional ASV system using the VoxCeleb1&2 datasets, without considering spoof robustness. We use a pre-trained model from ESPnet-SPK [60]. The second model (“SASV trained on out-of-domain data”) is trained as an SASV model but uses out-of-domain data from the ASVspoof2019 logical access dataset [9]. We use a pre-trained model from [61]. The third model (“SASV trained on SpoofCeleb”) is trained as an SASV model using the training set from SpoofCeleb.
@@ 6 Metrics
A diverse set of metrics is employed to evaluate the SpoofCeleb dataset, as well as the SDD and SASV models. To assess the quality of the speech samples and the strength of the attacks, we use SPF-EER, Mean Cepstral Distortion (MCD), UTMOS [62], DNSMOS [63], and Word Error Rate (WER), with the WER evaluated using the OpenAI Whisper-Large model [64]. SPF-EER measures speaker characteristics, UTMOS and DNSMOS are objective approximations of perceived quality and noisiness of synthesized speech, and WER measures intelligibility. For evaluating the performances of the SDD baselines, we use Equal Error Rate (EER) and the min Detection Cost Function (minDCF) [65]. To assess the SASV baselines, we adopt the recently proposed architecture-agnostic Detection Cost Function (min a-DCF) [66], along with Speaker Verification EER (SV-EER) and SPooF EER (SPF-EER). Table 5 outlines the trial types involved in the SASV metrics; a-DCF includes all three trial types, while SV-EER and SPF-EER cover only a subset.
Table 5: Three metrics used for gauging performances of SASV baselines. a-DCF measures the overall performance. SV-EER measures the ability to reject non-target speakers and SPF-EER measures spoof-robustness. “+”: a system should accept, “-”: a system should reject.
@@
Table 6: SDD baseline performances.
@@
Table 7: Attack-wise performance of RawNet2 SDD baseline on validation and evaluation sets. Performances reported using EER (%).
@@
Table 8: SASV baseline performances. SKA-TDNN [58] model architecture is employed. Three models are trained in different scenarios. “Conventional ASV” is trained with VoxCelebs1&2 [14, 67] and “SASV trained on out-of-domain data” is trained on ASVspoof2019 logical access [9].
@@ 7.1 Spoofing attacks
In the top row, the speech quality evaluations for A00 (bona fide speech) are provided as reference values. The results confirm that the spoofing attacks in SpoofCeleb are highly threatening, with most attacks achieving an SPF-EER over 20%. The majority of attacks exhibit relatively minor degradation in UTMOS and DNSMOS, indicating the high quality of the synthesized speech samples. Intelligibility, measured using the WER, shows that for most attacks, there is no more than a 10% deterioration in performance.
@@ 7.2 SDD
Table 6 presents the results of four baseline SDD systems. We evaluate two SDD models, RawNet2 and AASIST, trained on two different datasets. The models trained on the ASVspoof2019 logical access dataset are used to assess the zero-shot performance on validation and evaluation SDD protocols of SpoofCeleb. The other two models demonstrate the performance of systems trained on in-domain SpoofCeleb training data.
The zero-shot results in the top two rows indicate that existing SDD models not trained on in-the-wild data struggle to distinguish between spoofed samples and bona fide speech. As shown in rows 3 and 4, there is a significant performance improvement when these models are trained using the SpoofCeleb training set, highlighting the importance of training SDD models on in-the-wild data. However, the RawNet2’s result in row 3 is unexpected, as it shows better performance on the evaluation set than on the validation set, while the evaluation set includes totally unknown attacks. To further investigate this, we conduct an analysis of the attack-wise results.
@@
We found that the inferior performance on attack A11 contributed to the validation set results being worse than those on the evaluation set. Interestingly, when comparing A11 and A15, attack A15 is more difficult to distinguish for a conventional ASV system that does not account for spoofing, with SPF-EER values of 47.78% for A11 and 65.21% for A15. Both attacks originate from MQTTS; however, A11 was trained entirely from scratch, while A15 utilized a pre-trained decoder. Once an SASV system is trained on the SpoofCeleb training data, A11 becomes more challenging to detect. A deeper investigation into the reasons behind this phenomenon is left for future work.
The comparative analysis in Tables 4 and 7 reveals a discrepancy between the rankings of attacks’ SPF-EER on the pre-trained ASV system trained with VoxCeleb and the rankings of attacks’ EER on the SDD system trained with SpoofCeleb. This divergence may be attributed to the differences in training data, whether the models were trained on SpoofCeleb. The discrepancy could be a result of the fundamental differences in the tasks themselves, as SDD and SASV systems are optimized for distinct objectives.
@@ 7.3 SASV
Table 8 presents the performances of three SASV baselines on the SpoofCeleb validation and evaluation protocols. Min a-DCF assesses the overall performance, while SV-EER and SPF-EER evaluate the systems’ ability to reject bona fide and spoof non-target trials, respectively.
As expected, a conventional ASV system that does not account for spoof attacks, shown in the first row, fails to reject synthesized speech samples, with an a-DCF exceeding 0.49 on both the validation and evaluation sets. However, it performs well at rejecting bona fide non-target trials. The results in the second row indicate an improvement in a-DCF for the validation set, but even worse performance on the evaluation set. Both SV-EER and SPF-EER remain very high, indicating that the system trained for SASV with out-of-domain data struggles to reject both types of non-target trials. The a-DCF of 0.9998 also signifies that the model fails to find an operating point where it can reject both types of non-target trials. Finally, when trained on the SpoofCeleb training data, the a-DCF on the evaluation set drops to its lowest value (0.2902), and both SV-EER and SPF-EER are more balanced compared with row 1, where the system was only capable of rejecting bona fide non-target trials.